Citizens of the Nation-of-Blessed (NOB) on average consume a representative basket of goods and services annually. The value of the goods and services in the basket was Ghs 4,000 in the year 2012. In years 2013, 2014 and 2015 the cost of the same representative basket of goods and services, valued at the prices prevailing in the corresponding years, was Ghs5,000 , Ghs 6000, and Ghs 6,600 respectively.

a) Based on the above information, construct the consumer price index (CPI) series for NoB using year 2012 as the base or reference year.( Hint:note that the CPI=100.0 IN 2012).

b) What is the annual inflation rate, as measured by percentage in the CPI, in NoB for the years 2013, 2014, and 2015?

c)
1
Expert's answer
2016-11-24T12:06:14-0500
The value of the goods and services in the basket was:
Ghs 4,000 in 2012, Ghs 5,000 in 2013, Ghs 6,000 in 2014 and Ghs 6,600 in 2015.
a) The consumer price index (CPI) series for NoB using year 2012 as the base or reference year is:
CPI(2013) = 5,000/4,000*100 = 125,
CPI(2014) = 6,000/4,000*100 = 150,
CPI(2015) = 6,600/4,000*100 = 165.
b) The annual inflation rate, as measured by percentage in the CPI, in NoB for the years 2013, 2014, and 2015 is:
ri(2013) = (CPI(2013) - CPI(2012))/CPI(2012)*100% = (125 - 100)/100*100% = 25%,
ri(2014) = (CPI(2014) - CPI(2013))/CPI(2013)*100% = (150 - 125)/125*100% = 20%,
ri(2015) = (CPI(2015) - CPI(2014))/CPI(2014)*100% = (165 - 150)/150*100% = 10%.